# StickerConv: Generating Multimodal Empathetic Responses from Scratch

### 20 Jan 2024 v1

@@ 3.3 Plan Module
The Plan module is a core component of the Agent4SC, responsible for orchestrating the behavior of agents in empathetic interactions. To ensure that agents mimic human behavior in using stickers, we have specially designed three key actions to ensure the timely use of stickers, their coherence with the context, and their effective role in empathy.
  • Intention: Under the control of Profile module, the agent assesses the potential for using emotive stickers based on interaction history and observational data. The prompt template for this action is shown in Figure 11.
  • Query: If the agent intends to use a sticker, it will describe the emotional content of the sticker in detail. This description will be influenced by the Profile module, observation, and the specific intent behind using the sticker. The prompt template for this action is shown in Figure 12.
  • Select: The agent sends the query generated in the previous step to the sticker database accessible to it, retrieving the top-K stickers with the highest similarity. Subsequently, the agent combines the candidated sticker’s index, description, emotion and recommendation into a list. From this list, the agent selects the sticker it wishes to use. The prompt template for this action is shown in Figure 13.
3.4 Memory Module
In empathetic interactions, the coherence of communication and the consistency of personality traits are fundamental to the quality of responses. This necessitates that agents not only learn and provide feedback through observation during the interaction process, but also maintain long-term consistency in their personality traits. We adopted a strategy similar to that of Generative Agent Park et al. (2023): for short-term memory, we update the agent’s core traits based on the profile and observations; for long-term memory, the agent combines continuous interaction with external entities and regular self-reflection to maintain consistency.
@@ 3.5 Action Module
  • Message: generates text responses based on the Agent’s Profile and Memory module.
  • Sticker: engages in sticker usage as dictated by the Plan module.
@@ 3.6 Manager Agent
The Manager Agent, a critical componet in enhancing the performance of the Agent4SC, plays a significant role in ensuring interaction quality and dialogue consistency. This agent comprises two main modules:
Quality Assurance
This module is crucial for upholding high dialogue standards, It accomplishes this by: (1) Evaluating the relevance of stickers within the conversation, ensuring that they contribute meaningfully to the dialogue. (2) Regulating the use of stickers to prevent their overuse, thus maintaining the integrity and flow of the conversation. (3) Performing regular checks on dialogue quality, with a special focus on the appropriateness of response length and the avoidance of redundant information.
Process Control
This module takes charge of guiding the dialogue’s progression. It leverages historical interactions as a reference to make informed decisions on the optimal timing for concluding dialogues.
Figure 3: The chart of emotional distribution in the choice of stickers between users and the system. This chart revealed a striking trend: users have a significant preference for stickers that convey negative emotions, in contrast to the system. The system predominantly utilizes stickers expressing neutral and positive emotions. This comparison not only reflects the distinct emotional expression preferences between users and the system but also highlights the system’s active and supportive role in interactions.
@@ 3.7 The StickerConv Dataset
The Agent4SC facilitates empathetic interaction with humans through the collaborative interplay of its modules, and possesses the capability to utilize stickers. Leveraging this, we propose the construction of a large-scale multimodal empathetic dialogue dataset, StickerConv, with stickers serving as a non-textual modality.
Figure 4: The overview architecture of PEGS. The input images are jointly encoded by image encoder, Q-Former and a linear layer. Vicuna is used as the language model. To empower the language model to generate images, the output of the LLM is first mapped into the input space of the image decoder through a feature mapper, and then the frozen image decoder is employed to generate images.
In the dataset creation process, two roles are designated: User and System. The User, acting as the dialogue initiator, is assigned a profile randomly selected from 2,000 personality profiles generated as described in Section 3.1. The System, in the roles of a listener and empathizer, is assigned a predefined positive profile. Regarding the sticker tool, we have processed and subdivided the sticker dataset into 100 smaller vector databases. In each dialogue session, both User and System have access to one randomly assigned vector database from this pool. This arrangement is designed to better simulate human sticker usage behavior while enhancing the diversity and richness of stickers. It is important to note that vector databases numbered 1 to 80 are exclusively employed for the creation of training and cross-validation sets, whereas those numbered 81 to 100 are reserved for the construction of the test set.
StickerConv consists of 12,931 dialogue sessions, 67,505 stickers (unique 5.8K stickers) and over 2K conversation topic. The detailed statistical information of our dataset is shown in Table 4. On average, each dialogue session in StickerConv has 5.22 stickers and 5.49 turns. The emotional label distribution of the stickers used by User and System is shown in Figure 3. This analysis reveals notable differences in the emotional distribution of sticker usage, attributable to the distinct roles of the User and the System.
@@ 4 PEGS
We devise an empathetic large multimodal dialog model, PEGS, with the ability to PErceive and Generate Stickers. Figure 4 illustrates the overall structure of our model. Technically, we utilize ViT-G/14 from EVA-CLIP (Fang et al., 2023), Q-Former from BLIP-2 $\text{OPT}_{\text{2.7B}}$ (Li et al., 2023b), and a linear layer to encode images. Vicuna (Chiang et al., 2023), a widely used language model in LMMs, is employed for language modeling. We adopt Stable Diffusion (SD) (Rombach et al., 2022) as our image decoder.
@@ 4.1 Multimodal Input Perception
Multimodal Encoding
With reference to existing works (Li et al., 2023b; Dai et al., 2023; Zhu et al., 2023), we convert the multimodal inputs into feature vectors that can be solved by LLM. Specially, each text token is embedded into a vector $e_{t}\in\mathbb{R}^{1\times d}$, while each image is first encoded by a pre-trained vision encoder, and then a aligned feature vector $e_{v}\in\mathbb{R}^{32\times d}$ is obtained via Q-Former and a linear projection layer.
Training
Our effort is built on BLIP-2 (Li et al., 2023b), which has captured vision-language knowledge from numerous aligned image-text pairs. Zhu et al. (2023) have proven the effectiveness that freezing the pre-trained vision encoder, Q-Former, and LLM, while only pre-training a linear projection layer. As shown in Figure 4(a), we train the input linear projection layer using LAION115M, a sub-dataset of LAION400M (Schuhmann et al., 2021). The pre-training process undergoes 80,000 steps with a batch size of 32.
@@ 4.2 Multimodal Output Generation
Given original LLM vocabulary $V$ of size $|V|$, to enable the LLM to determine the timing of generating visual responses and to obtain feature vectors for image generation, we extend the vocabulary with an additional tokens set $V_{\mathrm{img}}=\{\mathrm{[IMG1]},\mathrm{[IMG2]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$. Therefore, the extended vocabulary $V^{*}=V\cup V_{\mathrm{img}}$. We denote the original word embedding matrix as $E\in\mathbb{R}^{|V|\times d}$. For the embeddings matrix $E^{*}\in\mathbb{R}^{|V^{*}|\times d}$ of the extended vocabulary $V^{*}$, the embeddings $E_{\mathrm{img}}\in\mathbb{R}^{r\times d}$ of the added special token $V_{\mathrm{img}}$ are randomly initialized and the embeddings $E$ of the original token are preserved:
$\centering E^{*}[0:|V|,:]=E\@add@centering$
@@
$\centering\mathcal{L}_{\mathrm{lm}}=-\sum_{i=1}^{k}\log p(u_{i}|s,u_{1},\ldots,u_{i-1};\theta,E_{\mathrm{img}})\@add@centering$
@@
Inspired by GILL (Koh et al., 2023), we align the hidden states $h_{\mathrm{gen}}$ corresponding to the output visual tokens $\{\mathrm{[IMG\{}i{\}]}\}_{i=1}^{r}$ into the input space of the image decoder. Specifically, we connect them through a feature mapper module, containing two linear layers and a 4-layer encoder-decoder transformer model with a learnable queries feature $q$. For the given image caption $c$ (and its emotion $e$ if available), our target is to minimize the MSE loss between their embeddings derived from the frozen pre-trained SD text encoder $\tau$ and the projected representations:
$\centering\mathcal{L}_{\mathrm{gen}}=\lVert\theta_{\mathrm{Mapper}}(h_{\mathrm{gen}},q)-\tau(c,[e])\rVert_{2}^{2}\@add@centering$
Training
Distinguished from realistic images, stickers have abstract nature and rich emotional expressions. Generic text-to-image models perform well in various scenarios, but they are flawed in sticker generation. Constrained by limited sticker data, there are not enough image-text pairs to turn the base text-to-image model to the style of stickers, and to pre-train for multimodal generation. As a remedy, we adopt a variant of CuteYukiMix [^1] https://civitai.com/models/28169/cuteyukimixadorable-style, a SD model trained on cute cartoon data, as the pre-trained text-to-image model. In the pre-training phase, our objective is to align the representations with the text embeddings (Figure 4(b)), so we train our model on the captions from LAION115M. The pre-training process undergoes 50,000 steps with a batch size of 32.
However, pre-trained SDs (focusing on obvious features, such as characters’ bodies, costumes) are insensitive to sentiment conditions, and it is difficult to directly perform end-to-end instruction tuning to learn emotion elements, we inserted a supervise fine-tuning (SFT) procedure on the sticker-text pairs with emotion to efficiently cause them to focus on emotion elements.
@@ 4.3 Joint Learning
Pre-trained LLMs have powerful text interaction capability, but perform poorly on open-domain conversations. We further utilize the constructed StickerConv for jointly fine-tuning of the entire model to achieve the capabilities of open-domain empathetic multimodal perception and generation. The instruction prompts we used are shown in Appendix C.4. As shown in Figure 4(c), we train our model in an end-to-end manner, using LoRA (Hu et al., 2021) to synchronize the update of a limited number of parameters in the LLM with the input linear projection layer and the feature mapper. The overall loss function $\mathcal{L}$ can be represented as:
$\centering\mathcal{L}=\lambda_{1}\times\mathcal{L}_{\mathrm{lm}}+\lambda_{2}\times\mathcal{L}_{\mathrm{gen}}\@add@centering$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. For $\mathcal{L}_{\mathrm{gen}}$, one of its inputs — emotion $e$ is available.
@@ 5 Experiments
Models IP EX ER diff-IP diff-EX diff-ER dist-1 dist-2 dist-3
Vicuna-7B 0.1857 0.2405 0.9509 0.3681 0.9745 0.2651 0.7995 0.9783 0.9936
ChatGLM3-6B 0.1632 0.2699 0.9768 0.344 0.9993 0.2629 0.7854 0.9776 0.9939
PEGS 0.1723 0.4458 0.9724 0.3113 0.8721 0.244 0.8397 0.9888 0.9976
Table 1: Results of quality of text generate in PEGS and baseline.
5.1 Experimental Settings
Implementation Details
At each training stage, the AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate with the initial learning rate of 1e-4. The details of pre-training have been introduced in Section 4. Our model is fine-tuned on the entire StickerConv. It is noteworthy that only 3% of the model parameters are updated, and it takes about 6 hours on a single A800 GPU.
Baseline
We compare with the following baselines: (1) Vicuna-7B: An open-source chatbot trained by fine-tuning LLaMA on user-shared conversations collected from ShareGPT Chiang et al. (2023). We use the Vicuna-delta-v0 version to ensure consistency with the model base used during training. (2) ChatGLM3-6B: An open-source, bilingual supported large-scale dialogue model Zeng et al. (2023).
5.2 Evaluation Metrics
5.2.1 General Metrics
This category focuses on assessing the quality of the model’s textual output.
Diversity
We adopt Distinct-n (Dist-1, 2, 3) Li et al. (2016) as quantitative indicators of diversity, measuring the variety and creativity of the model’s outputs.
Empathy Performance
To evaluate the model’s capacity for empathy, we adopt the EPITOME Sharma et al. (2020a) method, comprising three sub-metrics: IP (interpretations), EX (explorations) and ER (emotional reactions).
Empathy Difference
DIFF-EPITOME Lee et al. (2022) measures the difference scores of IP, EX, ER between the golden response and predicted response.
5.2.2 LLM-based Metrics
LLM has been widely proven to be capable of acting as a grader similar to humans and can provide reasonable scores based on the designed prompts. These metrics not only focues on textual outputs but also include sticker outputs, forming a comprehensive multimodal evaluation system.
Empathy
Utilizing ChatGPT-3.5-turbo as a scorer, based on the context, we assess empathy scores for the model’s textual and sticker responses, denoted as Emp-txt and Emp-sticker, respectively. To mitigate randomness and bias, we conduct five independent scorings for each response and calculate the average.
Consistency
Again using ChatGPT-3.5-turbo as a scorer, based on the context, we assign consistency scores for textual and sticker responses, marked as Con-txt and Con-sticker, employing the same scoring method.
Ranking
Employing ChatGPT-3.5-turbo as a result ranker, we compare responses of different models to the same context. The ranker organizes responses based on quality, empathy, and consistency, averaging across all possible ranking combinations to ensure fairness and objectivity in assessment.
5.2.3 Multimodal Metrics
CLIP-Score
We utilize CLIP (Radford et al., 2021) as multimodal encoder, evaluate the consistency of model’s multimodal responses.
MM-Relevance
MM-Relevance Feng et al. (2023) is a metric designed to measure the relevance of multimodal responses. Traditional mm-relevance metrics, however, face limitations, particularly in cases where there is a lack of alignment between model responses and golden responses. To address these challenges, we propose Inter-MM-Rel (Interactive multimodal Relevance). This innovative approach calculates and averages the similarity scores between each mode of information, without the constraint of alignment, allowing for a more comprehensive and accurate assessment of multimodal response relevance.

### 16 Feb 2024 v2

@@ 3.4 Plan Module
The Plan module orchestrates agent behavior in empathetic interactions. To ensure that agents mimic human behavior in using stickers, we specially design three key actions to ensure the timely use of stickers, their coherence with the context, and their effective role in empathy.
  • Intention: Directed by the Profile module, the agent assesses the use of emotive stickers, considering interaction history and observed data.
  • Query: When intending to use a sticker, the agent describes its emotion, shaped by the Profile module, observations, and the intent for its use.
  • Select: After querying its sticker database with the crafted query, the agent retrieves the top-K most relevant stickers. It organizes these stickers into a list detailing their emotion, description, and usage recommendation, from which it selects the most appropriate sticker.
@@ 3.5 Action Module
  • Message: The agent generates text responses based on the agent’s Profile and Memory module.
  • Sticker: The agent engages in sticker usage as dictated by the Plan module.
@@ 3.6 Manager Agent
The Manager Agent, a critical component in enhancing the performance of the Agent4SC, plays a significant role in ensuring interaction quality and dialogue consistency. This agent comprises two main modules: (1) Quality Assurance: This module plays a key role in maintaining high dialogue standards by evaluating sticker relevance to enrich conversations, regulating sticker usage to avoid misuse, and conducting quality checks focused on response length appropriateness and redundancy avoidance. (2) Process Control: This module takes charge of guiding the dialogue’s progression. It leverages historical interactions as a reference to make informed decisions on the optimal timing for concluding dialogues.
@@ 3.7 The STICKERCONV Dataset
Leveraging Agent4SC, we propose the construction of a multimodal empathetic dialogue dataset, STICKERCONV.
Figure 3: The architecture of PEGS framework includes various routing options, distinguished by colored connecting lines. Input stickers undergo joint encoding by an image encoder, Q-Former, and a linear layer, with Vicuna serving as the language model. The output of the LLM activates two sets of tokens differently across model versions: one for image retrieval and the other as a textual condition. Subsequently, the frozen image decoder generates images.
In our dataset creation, we define two roles: User, the dialogue initiator with a profile from 2,000 generated personalities, and System, acting as a listener and empathizer. To mimic the human use of stickers, the sticker dataset is divided into 100 vector databases, each ensuring a consistent emotional distribution of stickers. The top-K of Tool module is set to 10. For each dialogue session, both User and System access a randomly chosen vector database, mirroring human sticker usage and enriching sticker diversity.
STICKERCONV consists of 12,931 dialogue sessions, 67,505 stickers (unique 5.8K stickers) and 2K user personalities. Each session averages 5.22 stickers and 5.49 dialogue turns. Emotional label distribution analysis, as illustrated in Figure 4, highlights the differences in sticker usage between User and System, reflecting their unique roles. Further analysis of the STICKERCONV is detailed in the Appendix A.
@@ 4 PEGS
We devise a multimodal empathetic response generation framework, PEGS, with the ability to PErceive and Generate Stickers. Figure 3 illustrates the architecture of our framework. With different strategies to generate images, we derive three models based on this framework: PEGS-Ret/Gen/RAG, which denote the retrieval, generation, and retrieval-augmented generation methods to provide images. Technically, we utilize ViT-G/14 from EVA-CLIP (Fang et al., 2023), Q-Former from BLIP-2 (Li et al., 2023b), and a linear layer to encode images. Vicuna (Chiang et al., 2023), a widely used language model in LMMs, is employed for language modeling. Stable Diffusion (SD) (Rombach et al., 2022) as the image decoder.
@@ 4.1 Multimodal Input Perception
With reference to existing works (Li et al., 2023b; Dai et al., 2023; Zhu et al., 2023), we convert the multimodal inputs into feature vectors that can be solved by LLM. Specially, each text token is embedded into a vector $e_{t}\in\mathbb{R}^{1\times d}$, while each image is first encoded by a pre-trained vision encoder, and then an aligned feature vector $e_{v}\in\mathbb{R}^{32\times d}$ is obtained via Q-Former and a linear projection layer.
@@ 4.2 Multimodal Output Generation
We extend the vocabulary $V$ with an additional visual tokens set $V_{\mathrm{img}}=\{\mathrm{[IMG1]},\mathrm{[IMG2]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$. We denote the original word embedding matrix as $E\in\mathbb{R}^{|V|\times d}$. For the embeddings matrix $E^{*}\in\mathbb{R}^{|V^{*}|\times d}$ of the extended vocabulary $V^{*}=V\cup V_{\mathrm{img}}$, the embeddings $E_{\mathrm{img}}\in\mathbb{R}^{r\times d}$ of the added special tokens are randomly initialized and the embeddings $E$ of the original token are preserved:
$\centering E^{*}[0:|V|,:]=E\@add@centering$ (1)
We split visual tokens into two sets, where the front $k$ tokens for image retrieval and the back $r-k$ tokens for image generation:
$V_{\mathrm{ret}}=\{\mathrm{[IMG1]},\ldots,\mathrm{[IMG\{}k\mathrm{\}]}\}$ (2)
$V_{\mathrm{gen}}=\{\mathrm{[IMG\{}k+1\mathrm{\}]},\ldots,\mathrm{[IMG\{}r\mathrm{\}]}\}$ (3)
where $V_{\mathrm{ret}}$ is used in PEGS-Ret and PEGS-RAG and $V_{\mathrm{gen}}$ is used in PEGS-Gen and PEGS-RAG.
@@
$\centering\mathcal{L}_{\mathrm{lm}}\!=\!-\!\sum_{i=1}^{k}\!\log p(u_{i}|s,u_{1},\ldots,u_{i-1};\theta,E_{\mathrm{img}})\@add@centering$ (4)
@@
Image Retrieval
For image retrieval, PEGS aligns the hidden states $h_{\mathrm{ret}}$ corresponding to $V_{\mathrm{ret}}$ into the retrieval space by contrastive learning (Chopra et al., 2005). ${W}_{\mathrm{t}}\in\mathbb{R}^{d\times e}$ and ${W}_{\mathrm{i}}\in\mathbb{R}^{p\times e}$to bridge the semantic gap and adjust the dimension. Cosine similarity is used to measure the similarity of the projection vectors:
$\centering\text{sim}(x,y)=\frac{({W}_{\mathrm{t}}^{T}h_{\mathrm{ret}}(x))^{T}({W}_{\mathrm{i}}^{T}\nu_{\phi}(y))}{\|{W}_{\mathrm{t}}^{T}h_{\mathrm{ret}}(x)\|\|{W}_{\mathrm{i}}^{T}\nu_{\phi}(y))\|}\@add@centering$ (5)
where $\nu_{\phi}$ is the image encoder. The projection vectors are used to minimize the InfoNCE loss (Oord et al., 2018), which consists of text-to-image (t2i) and image-to-text (i2t) loss in a batch of N text-image pairs ($x_{i}$, $y_{i}$):
$\mathcal{L}_{\mathrm{t2i}}\!$ $=\!-\frac{1}{N}\!\sum_{i=1}^{N}\!\left(\!\log\!\frac{\exp(\text{sim}(x_{i},y_{i})\!/\!\tau)}{\sum_{j=1}^{N}\!\exp(\text{sim}(x_{i},y_{j})\!/\!\tau)}\!\right)$ (6)
$\mathcal{L}_{\mathrm{i2t}}\!$ $=\!-\frac{1}{N}\!\sum_{i=1}^{N}\!\left(\!\log\!\frac{\exp(\text{sim}(x_{i},y_{i})\!/\!\tau)}{\sum_{j=1}^{N}\!\exp(\text{sim}(x_{j},y_{i})\!/\!\tau)}\!\right)$ (7)
$\centering\mathcal{L}_{\mathrm{ret}}=\frac{1}{2}(\mathcal{L}_{\mathrm{t2i}}+\mathcal{L}_{\mathrm{i2t}})\@add@centering$ (8)
$\mathcal{L}_{\mathrm{ret}}$ is the loss used to optimize the projection layers for retrieval.
@@
We align the hidden states $h_{\mathrm{gen}}$ corresponding to the output visual tokens $V_{\mathrm{gen}}$ into the input space of the image decoder. Specifically, we connect them through a feature mapper module, containing two linear layers and a 4-layer encoder-decoder transformer model with a learnable queries feature $q$. For the given image caption $c$ (and its emotion $e$ if available), our target is to minimize the MSE loss between their embeddings derived from the frozen pre-trained SD text encoder $\eta$ and the projected representations:
$\centering\mathcal{L}_{\mathrm{gen}}=\lVert\theta_{\mathrm{Mapper}}(h_{\mathrm{gen}},q)-\eta(c,[e])\rVert_{2}^{2}\@add@centering$ (9)
Retrieval-Augmented Image Generation
Intuitively, continuing to do generation on retrieved images can extend the diversity of images while maintaining image quality, thus we explore retrieval-augmented generation.
Specifically, we retrieve an image serving as a latent representation $c_{I}$ for augmenting the generation process. During image generation, $h_{\mathrm{gen}}$ remains utilized as a condition.
@@ 4.3 Joint Learning
Pre-trained LLMs excel in text interactions yet struggle in empathetic conversations. We further utilize the constructed STICKERCONV for jointly fine-tuning of the entire model to achieve the capabilities of multimodal perception and generation. We train our model in an end-to-end manner, using LoRA (Hu et al., 2021) to synchronize the update of a limited number of parameters in the LLM with the input linear projection layers and the feature mappers. The overall loss function $\mathcal{L}$ can be represented as:
$\centering\mathcal{L}=\lambda_{1}\times\mathcal{L}_{\mathrm{lm}}+\lambda_{2}\times\mathcal{L}_{\mathrm{gen}}+\lambda_{3}\times\mathcal{L}_{\mathrm{ret}}\@add@centering$ (10)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ represent hyperparameters. $\mathcal{L}_{\mathrm{gen}}$ includes emotion $e$ as an input. For PEGS-Ret, $\mathcal{L}_{\mathrm{gen}}=0$, and for PEGS-Gen, $\mathcal{L}_{\mathrm{ret}}=0$. Implementation details can be found in Appendix C.
5 Evaluation Metrics
5.1 Text Metrics
To comprehensively evaluate the fluency, diversity and accuracy of dialogue generation, we utilize an array of broadly recognized text metrics, comprising BLEU Papineni et al. (2002), Distinct-n (Dist-n) Li et al. (2016) , ROUGE-L (ROU_L) Kingma and Ba (2017), METEOR (MET) Banerjee and Lavie (2005), CIDEr Vedantam et al. (2015), BERTScore (BERTS) Zhang et al. (2020).
5.2 Multimodal Metrics
We use MM-Relevance (MMr) Feng et al. (2023) to assess the relevance between the predicted multimodal response and the golden response. However, MMr overlooks the frequency of multimodal replies, a critical aspect considering that text responses within a modality often show higher similarity than across modalities. This could result in models that rarely engage in multimodal responses receiving inaccurately high MMr scores. To address this issue, we impose penalties on models that are more inclined to generate text-only replies:
$f\text{-MMr.}=(1-\alpha(1-f))\cdot\text{MMr.}$. (11)
where $\alpha$ represents the penalty coefficient. In our experiments, we set $\alpha$ to 0.8. $f$ denotes the relative frequency at which the model produces multimodal responses, $f\in[0,1]$.
Model BLEU-1/2/3/4 Dist-1/2/3 ROU_L. MET. CIDEr BERTS.
Vicuna-text 0.44/0.30/0.22/0.17 0.879/0.994/0.999 0.31 0.37 0.39 0.878
Vicuna-tool 0.43/0.29/0.22/0.17 0.870/0.989/0.994 0.30 0.36 0.38 0.900
ChatGLM3-text 0.42/0.28/0.21/0.16 0.806/0.981/0.996 0.31 0.40 0.40 0.886
ChatGLM3-tool 0.36/0.22/0.16/0.11 0.859/0.992/0.998 0.26 0.34 0.20 0.899
PEGS-Ret 0.46/0.32/0.25/0.20 0.839/0.989/0.997 0.34 0.42 0.47 0.906
PEGS-RAG 0.46/0.32/0.25/0.20 0.839/0.989/0.997 0.34 0.42 0.47 0.906
PEGS-Gen 0.47/0.33/0.26/0.21 0.848/0.990/0.997 0.35 0.44 0.57 0.911
Table 1: Results of quality of text generate in PEGS and baseline models.
5.3 LLM-based Metrics
LLMs are capable of grading similarly to humans, providing scores for both textual and sticker outputs, thus enabling a comprehensive multimodal evaluation system.
We introduce three LLM-based metrics: (1) Empathy: We assess empathy in model responses, both textual (Empathy-text, EMP-txt) and multimodal (Empathy-multimodal, EMP-mm), averaging scores from five independent scorings to reduce randomness and bias. (2) Consistency: Based on the context, we assign consistency scores for textual and sticker responses, marked as Consistency (CON), employing the same scoring method. (3) Rank: We compare responses of different models to the same context. The ranker organizes responses based on quality, empathy, and consistency, averaging across many possible ranking combinations to ensure fairness and objectivity.
5.4 Human Metrics
Considering the subjectivity of empathy and the complexity of multimodal responses, we establish seven detailed manual evaluation metrics: Sticker Generation Quality (StiGQ), Empathy Sticker (Es), Empathy Text (Et), Consistency (Con), Fluency (Flu), Informativity (Inf), Sticker Diversity (StiD).
Metrics are scored on a 1 to 5 scale, with higher scores denoting better performance. For fairness and consistency, all models are evaluated in identical contexts. The evaluation panel consists of five members who score models anonymously, unaware if they are assessing a baseline model. The human evaluation process randomly selects 100 dialogue sessions from the STICKERCONV test set.
@@ 6 Experiments
6.1 Implementation Details
Our models are pre-trained on LAION115M, a sub-dataset of LAION400M (Schuhmann et al., 2021), and fine-tuned on STICKERCONV. More implementation details and hyperparameter settings are provided in Appendix C. The training procedure is conducted on 2 NVIDIA A6000 48G GPUs.
6.2 Baselines
We benchmark against Vicuna-7B Chiang et al. (2023) and ChatGLM3-6B Zeng et al. (2023) models, employing two experimental paradigms:
Text Fine-tuning
We fine-tune the models using text data from the STICKERCONV training set and then make predictions on the test dataset.
Tool Learning
We fine-tune the models with multimodal data from the STICKERCONV training set. We use stable-diffusion as a tool to teach the models how to utilize it for generating sticker responses Tang et al. (2023). To ensure fairness, the tool models employ the same SD model as PEGS.